OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3? y = –x – 1 y = –x + 5 y = x + 1 y = x + 5

OpenStudy (anonymous):

how do i do this

OpenStudy (anonymous):

First thing, do you know the slope of y=-x-3?

OpenStudy (anonymous):

no how do i get it

OpenStudy (anonymous):

its already in the form y=mx+b where m is the slope, so the coefficient of x is the slope

OpenStudy (anonymous):

so the slope is -1?

OpenStudy (anonymous):

Yep, a line perpendicular to that would have the negative reciprocal of its slope. so slope of the line you're trying to find is 1 use point slope form now y-y1=m(x-x1)

OpenStudy (anonymous):

wait what do you mean use point slope now ?

OpenStudy (anonymous):

@CalebBeavers

OpenStudy (anonymous):

point slope form y-y1=m(x-x1) (x1,y1) is the point you have (2,3) and m is the slope, 1 plug those in